Linda’s Luxury Travel (LLT) is considering the purchase of two Hummer limousines. Various information about the proposed investment follows:

Intital Investment 780,000
Useful Life 10 Years
Salvage Value 100,000
Annual Net Income Generated 64,740
LLT'S Cost of capital 15%

Assume straight line depreciation method is used.     

Required:

Help LLT evaluate this project by calculating each of the following:

1. Accounting rate of return.

2. Payback period.

3. Net present value.

4. Without making any calculations, determine whether the IRR is more or less than 15%.
Answer #1

Initial investment = 780000

Depreciation = 780000 - 100000/ 10 = 68000

1.accounting rate of return = average profits/ initial investment

= 64740 / 780000 = 8.3 %

Cash inflows = annual income + depreciation = 64740 + 68000

= 132740

2. Payback period = initial investment/ annual cash inflows

= 780000/ 132740 = 5 years 10.5 months

3.npv = present value of discounted cash inflows- initial investment

132740 × 4.771 = 633302.54

132740 + 100000 = 232740 × 0.247 = 57486.78

Present value of cash inflows = 690789.32

Npv = 690789.32 - 780000 = 89210.68 Negative

4. Irr is at which npv is equal to zero so, the irr will be less than 15%
Answer #2

Initial Investment = $780,000
Salvage Value = $100,000
Useful Life = 10 years

Annual Depreciation = (Initial Investment - Salvage Value) / Useful Life
Annual Depreciation = ($780,000 - $100,000) / 10
Annual Depreciation = $68,000

Annual Net Cash Flow = Annual Net Income + Annual Depreciation
Annual Net Cash Flow = $64,740 + $68,000
Annual Net Cash Flow = $132,740

Answer 1.

Accounting Rate of Return = Annual Net Income / Initial Investment
Accounting Rate of Return = $64,740 / $780,000
Accounting Rate of Return = 8.30%

Answer 2.

Payback Period = Initial Investment / Annual Cash Flow
Payback Period = $780,000 / $132,740
Payback Period = 5.88 years

Answer 3.

Cost of Capital = 15%

Net Present Value = -$780,000 + $132,740 * PVA of $1 (15%, 10) + $100,000 * PV of $1 (15%, 10)
Net Present Value = -$780,000 + $132,740 * 5.0188 + $100,000 * 0.2472
Net Present Value = -$89,084.48 or -$89,084

Answer 4.

Internal rate of return is less than 15% as NPV is negative at cost of capital of 15%.